he fundamental accounting equations for several businesses follow.

Supply the missing amounts. Assets = Liabilities + Owner’s Equity

Assets = Liabilities + Owner’s Equity
1. $27,100 = $5,210 +
2. $23,500 = $4,830 +
3. $48,375 = + $10,225
4. = $4,050 + $32,150
5. $54,300 = + $24,950
Answer #1
Assets = Liabilities + Owner’s Equity
1 27100 = 5210 + 21890 =27100-5210
2 23500 = 4830 + 18670 =23500-4830
3 48375 = 38150 + 10225 =48375-10225
4 36200 = 4050 + 32150 =4050+32150
5 54300 = 29350 + 24950 =54300-24950